OpenStudy (anonymous):

Find the derivative of the function. y=(3 x^2 - 6 x + 1)^5

OpenStudy (ash2326):

We have \[ y=(3 x^2 - 6 x + 1)^5 \] \[\frac{dy}{dx}= 5\times (3x^2-6x+1)^4 \times \frac{d}{dx}(3x^2-6x+1)\] we get \[\frac{dy}{dx}= 5\times (3x^2-6x+1)^4 \times (6x-6)\]

OpenStudy (campbell_st):

use the chain rule dy/dx = dy/du x du/dx dy/dx = 5(3x^2 - 6x +1)^4 x (6x - 6)

OpenStudy (anonymous):

ok is 5x? or 5 times (x)?

OpenStudy (ash2326):

5 times

OpenStudy (anonymous):

thanx